# Revisiting Few-Shot Object Detection with Vision-Language Models

### v2

@@ 4.5 Fine-Tuning Ablation
Detic Components 10 Shots: Average Precision (AP)
Backbone RPN Box Regressor Classifier All Many Medium Few
- - - - 12.11 19.41 18.44 0.87
[…]
4.6 Limitations and Future Work
Despite using VLMs pre-trained on large-scale datasets, we find that performance for rare categories (as defined by the cardinality of each class in the original dataset) is considerably lower than for common classes. We posit that VLMs are pre-trained with imbalanced data which includes many examples of common categories like truck but few examples of rare categories like stroller. Our work does not explicitly improve detection performance on the rare classes. Interestingly, since VLMs like Detic [49], GLIP [20], and GroundingDINO [23] are trained with different data sources, each model has dramatically different zero-shot performance on novel categories like stroller. Ensembling predictions from different VLMs may yield better detection accuracy for rare categories. In addition, although our work motivates the use of rich textual descriptions for multi-modal alignment, our approach only uses class names as text features. We hope future work can address the above limitations.
5 Conclusion
We revisit few-shot object detection (FSOD) with vision-language models (VLMs) and find that zero-shot inference from web-scale VLMs significantly outperforms leading FSOD methods. However, such foundational models do not fully address few shot recognition because of the concept alignment problem; particular concepts in target applications may be different than their use on web-scale datasets. Just as human annotators require concept alignment via multimodal text and visual examples, we argue that VLMs should be aligned with such few-shot data, formalizing the problem of Foundational FSOD. We also point out that existing FSOD benchmarks are actually federated datasets, and demonstrate that federated losses improve Foundational FSOD performance, approaching the oracle upper bound where few-shot images are fully annotated. Our analysis suggests that future few-shot (or “foundational concept alignment”) benchmarks may benefit from assuming images are fully annotated, since the cost of annotating a small set of $K$ images is manageable in practice.
Appendix A Implementation Details
Pseudo-Negative Federated Loss. First, we sample an image at random. We start by running a pre-trained detector on the image to generate predictions. We partition the predictions into two sets based on a confidence threshold; all predictions above the threshold are pseudo-positives (pseudo_pos) and all predictions below the threshold are ignored. Next, we compute negative classes (neg_classes) based on the pseudo-positives. Specifically, all classes not included in the psuedo-positive predictions are considered negative classes. Following the federated loss described in [48] (modified to use neg_classes instead of randomly sampled negative classes), we compute the binary cross entropy loss for the negative (neg_classes) and ground truth classes. See Algorithm 1 for pseudo-code.
# Inputs
@@ Appendix A Implementation Details
LVIS v0.5 Experiment Details. We select Detic with a Resnet-50 backbone for fair comparison with prior work. We pre-train Detic on LVIS-base for $90k$ iterations with a batch size of $32$ using an AdamW optimizer and a learning rate of $32e-$ 3. All images are resized to $640\times 640$ and we also enable Repeat Factor Sampling [11].
Following [38], we sample up to $10$ shots for each class in LVIS (since all classes may not have 10 examples). We use a batch size of $32$, learning rate of $52.5e-$ 5 for $46k$ iterations. We do not use Repeat Factor Sampling for fine-tuning. We sample $50$ categories for each training image, i.e $|S|=50$ for the FedLoss and InvFedLoss experiments. We derive negatives from pseudolabels with atleast $20\%$ confidence for the Psuedo-Negative experiment.
nuImages Experiment Details. We select Detic with a Swin-T backbone pre-trained on LVIS+COCO and ImageNet-21k data. We use an image size of $1600\times 900$, batch size of $8$ and an AdamW optimizer with learning rate of $63.75e-$ 6. We fine-tune this model for $8000$ iterations on nuImages. We sample $6$ categories for each training image, i.e $|S|=6$ for the FedLoss and InvFedLoss experiments. We derive negatives from pseudolabels with atleast $20\%$ confidence for the Psuedo-Negative experiment.
The original Detic implementation samples $|S|$ categories per batch, rather than per image. This works for LVIS, which has 1200 categories. However, nuImages only has $18$ classes, which required re-implementing the negative category sampling step for each image.
Appendix B Analysis of Iconic Few-Shot Images
The specific examples used during few-shot fine-tuning significantly impacts target class performance [38]. However, prior work constructs few-shot splits by randomly sampling $K$ examples per class. In contrast, when creating annotator instructions, selecting the right examples to “align” human annotators [2] to subtle aspects of the target concept is carefully considered. To more closely match VLM concept alignment with human annotator alignment, we design a simple algorithm to construct the best $K$-shot split for fine-tuning. This allows us to understand which examples are most informative and measure an upper bound in performance.
We construct our best split by picking examples corresponding to the best class-wise performance, based on the evaluation of each split on a held-out validation set. For instance, out of $10$ random splits for the $5$-shot task, one may pick car examples from split $1$, bicycle from split $4$ and debris from split $8$. In Table 6, we observe that the best-split performance is always better than its random counterpart. As expected, the choice of examples in $5$-shot case is more important than the $30$-shot case ($1.63$ AP difference for $5$-shot vs $0.08$ AP for $30$-shots). We visualize the difference in the splits for the stroller class in nuImages (cf. Figure 5). Unsurprisingly, iconic examples are large and unoccluded.
Figure 5: Visualizing Random and Best Split. In the top row, we visualize the 5-shot training examples of strollers from a random split. Similarly, we visualize the 5-shot training examples from the best split in the bottom row. We observe that strollers in the random split are often occluded, small in size and are blurry, making few-shot learning harder. On the other hand, the best split examples are larger, have better visual quality and are relatively un-occluded. This visual difference directly translates into better few-shot performance. We achieve $\mathbf{13.09}$ Stroller AP for the random split and $\mathbf{18.54}$ Stroller AP for the best split. We perform a more comprehensive evaluation in Table 6.
Table 6: Random Split vs ‘Best” Split. We construct the “best” split by selecting per-class few-shot examples that lead to the highest performance on a held out set. Unsurprisingly, the best split performs better than any random split, especially for very limited data settings (e.g. $5$-shot detection). This evaluation setting closely mimics how human annotators are “aligned” to target concepts, since annotator guides are constructed using hand-picked iconic visual examples.
Approach Average Precision (AP)
@@ Appendix B Analysis of Iconic Few-Shot Images
Detic (Zero-Shot) [49] 14.26 27.28 16.88 2.36
5-shots
   Ours (Random Split) 15.94 28.47 19.53 3.50
   Ours (Best Split) 17.57 29.99 20.70 4.92
10-shots
   Ours (Random Split) 16.67 29.15 20.66 3.90
   Ours (Best Split) 17.51 29.98 21.60 4.58
30-shots
   Ours (Random Split) 17.87 30.32 22.35 4.70
   Ours (Best Split) 17.95 29.50 22.69 4.85
Appendix C nuImages Foundational FSOD Performance for K={5, 30} Shots
We observe similar trends for $K=\{5,30\}$ to the 10-shot performance reported in the main paper. Note that the reported results are averaged over $10$ seeds to reduce variance. For $K=5$ shots, we note that naive fine-tuning barely improves the overall AP (cf. Table  7). Incorporating pseudo-negatives in this data-scarce setting closes the gap to the true negatives upper bound ($15.94$ vs $16.15$ AP). Even with exhaustive annotations, we only improve by $\sim 3.3$ AP ($14.26$ vs $17.59$ AP) points over the zero-shot results.
Naive fine-tuning improves over zero shot results by $\sim 2.5$ AP ($16.83$ vs $14.26$ AP) in the $K=30$ shot setting. In addition, we see diminishing benefits for using FedLoss variants (cf. Table 8). Nevertheless, pseudo-negatives improve over naive fine-tuning by $1$ AP ($16.83$ vs $17.87$ AP). Unsurprisingly, performance for classes with many examples increases the most ($19.40$ AP) when training with exhaustive annotations. Importantly, training with exhaustive annotations uses orders of magnitude more examples (1300 examples of car) than with few-shot annotations (30 examples of car).
Table 7: Analysis of $5$-shot Performance on nuImages. Fine-tuning with pseudo-negatives improves over naive fine-tuning. Importantly, our pseudo-negative approach nearly matches the upper bound with true negatives ($15.94$ AP vs $16.15$ AP). This suggests that our proposed approach is particularly effective in data-constrained settings.
Approach 5 shots: Average Precision (AP)
All Many Medium Few
Detic [49] Zero-Shot 14.26 27.28 16.88 2.36
   + Fine-tuning             14.77 25.16 18.65 3.63
   w/ FedLoss 15.03 26.41 19.07 3.03
   w/ Inverse FedLoss 15.21 26.77 19.01 3.29
   w/ Pseudo-Negatives 15.94 28.47 19.53 3.50
   w/ True Negatives (Oracle) 16.15 29.00 19.52 3.70
   w/ Exhaustive Annotations (Oracle) 17.59 32.66 20.93 3.43
Table 8: Analysis of $30$-shot performance on nuImages Naive fine-tuning provides a considerable improvement over zero-shot inference due to a greater number of examples per-class. In constrast to 5-shot results, FedLoss variants provide limited improvement over naive fine-tuning. We find that fine-tuning with pseudo-negatives provides a 1% improvement overall.
Approach 30 shots: Average Precision (AP)
All Many Medium Few
Detic [49] Zero-Shot 14.26 27.28 16.88 2.36
   + Fine-tuning             16.83 27.90 21.59 4.45
   w/ FedLoss 16.45 28.88 21.14 3.02
   w/ Inverse FedLoss 16.88 29.19 21.41 3.77
   w/ Pseudo-Negatives 17.87 30.32 22.35 4.70
   w/ True Negatives (Oracle) 18.19 30.61 22.66 5.11
   w/ Exhaustive Annotations (Oracle) 19.40 34.24 23.75 4.48
@@
RegionCLIP (Zero-Shot) [47] 2.34 3.33 3.87 0.22
Detic (Zero-Shot) [49] 14.26 27.28 16.88 2.36
5-shots
   RegionCLIP (Fine-Tuning) [47] 3.61 6.20 5.14 0.26
   Detic (Fine-Tuning) [49] 14.50 24.09 18.53 3.70
10-shots
   RegionCLIP (Fine-Tuning) [47] 3.58 6.10 5.16 0.24
   Detic (Fine-Tuning) [49] 15.28 26.93 19.89 3.27
30-shots
   RegionCLIP (Fine-Tuning) [47] 3.57 6.13 5.10 0.22
   Detic (Fine-Tuning) [49] 16.65 27.45 21.51 4.02
Appendix D RegionCLIP Experiments
In this section, we evaluate the importance of using box supervised data in pre-training. Unlike Detic, which trains on box-supervised data from LVIS, COCO and image-text data from ImageNet21k, RegionCLIP[47] only pre-trains on image-text pairs from the Conceptual Captions (CC3M) dataset [35].
Table 10: Diagnosing RegionCLIP’s Poor Zero-Shot Performance. RegionCLIP’s zero-shot performance lags far behind Detic. Using RegionCLIP’s classifier on ground-truth region proposals yields high performance, suggesting that RegionCLIP struggles to accuratly distinguish between foreground-vs-background.

### v3

@@ 5 Conclusion
Appendix A Baseline Implementation Details
We repurpose nuImages (CC BY-NC-SA 4.0) for all few-shot experiments in the main paper. We evaluate detection performance using $1600\times 900$ images across 18 classes for all models tested. We create three random splits for each of $K=\{5,10,30\}$-shots following the data creation process from [wang2020frustratingly] and report results averaged across these three seeds. Our test-set is a subset of the (densely annotated) nuImages val-set. We construct our test-set to only include validation images which have at least one annotation from the Few or Medium cohorts (cf. Fig 4). We train all baselines with one RTX 3090 GPU. Our baseline code is available on GitHub and dataset splits are available on HuggingFace.
Figure 4: We visualize the distribution of classes in out test-set compared to the cardinalities of classes in the full nuImages val-set. Notably, our sub-sampling strategy of selecting validation images that have at least one annotation from medium or few classes does not significantly alter the true distribution.
Prompt Engineering: We leverage rich text descriptions provided by the annotator instructions to select synonyms for each nuImages class. We manually identify the best performing synonyms in Table 5. At test time, we compute the average text embedding of all synonyms to improve classification accuracy.
Table 5: Synonyms used for Prompt Engineering. We manually inspect the nuImages annotator instructions to derive a set of synonyms to improve classification accuracy.
Original Classes Class Names with Synonyms
car car
truck truck, pick-up, lorry, semi-tractor
construction_vehicle construction_vehicle, crane
bus bus, bendy_bus, rigid_bus
trailer trailer
emergency emergency, ambulance, police_car, police_motorcycle
motorcycle motorcycle
bicycle bicycle
adult adult, person
child child
police_officer police_officer
construction_worker construction_worker
personal_mobility personal_mobility, skateboard, segway, scooter
stroller stroller
pushable_pullable pushable_pullable, wheel_barrow, garbage_bin, cart
barrier barrier, K-rail, fence, bollard, guard_rail
traffic_cone traffic_cone
debris debris, trash_bag
Language Prompt Tuning We train GLIP (SWIN-L backbone) for our prompt tuning experiments for $60$ epochs with a learning rate of $0.025$, batch size of $4$, and weight decay of $0.25$.
Federated Fine-tuning. We use Detic (Swin-B backbone) pre-trained on LVIS + COCO and ImageNet-21k data for our federated fine-tuning experiments (described in detail in the next section). We use a batch size of $8$ and an AdamW optimizer with learning rate of $63.75e-$ 6. We fine-tune this model for $8000$ iterations on nuImages. We sample $6$ categories for each training image, i.e $|S|=6$ for the FedLoss and InvFedLoss experiments. We derive negatives from pseudolabels with atleast $20\%$ confidence for the Psuedo-Negative experiment.
Multi-Modal Prompting. We use MQDet (text-only, vision-only, text + vision for our in-context learning baselines. Unlike the original code base, we tokenize our few shot examples instead of using random queries. Note that zero-shot results for MQ-GLIP-Text and GLIP-L are the same since these models are identical.
Appendix B CVPR 2024 Competition Details.
Table 6: CVPR 2024 Foundational FSOD Competition Results.
Team Name Average Precision (AP)
All Many Medium Few
PHP_hhh 45.35 64.25 53.43 20.19
NJUST KMG 32.56 50.21 34.87 15.16
zjyd_cxy_vision 31.57 46.59 33.32 17.03
Baseline (MQ-GLIP) 21.51 32.25 23.35 10.41
team_anon 17.36 25.29 21.93 5.42
youyouqiu 13.16 11.29 19.20 7.68
zhao 11.38 11.16 16.76 5.30
zjdcxy 7.80 5.44 13.43 3.20
Our inaugural Foundational FSOD competition (hosted on Eval AI) received submissions from eight teams (some submissions are private) around the world. We present a ranked list of participants at the close of our competition on June 7th AOE in Table 6. Notably, three teams were able to beat our MQ-GLIP baseline. Unfortunately, the top performing team wasn’t willing to publicly share details about their method. We summarize contributions from the other two top teams below.
NJUST KMG presents a method leveraging a vision-language model (VLM) enhanced with a multimodal large language model (MM-LLM) to improve Few-Shot Object Detection (FSOD). To address the challenge of misalignment between VLMs and target concepts, authors propose generating descriptive referential expressions for each category using MM-LLM. This involves annotating images with bounding boxes, prompting ChatGPT to provide descriptive terms for each object, and then creating multiple referential expressions by randomly combining these terms. The VLMs then select the best referential expression for each category by matching the maximum Intersection over Union (IoU) in the training set, and these expressions are used to generate pseudo-labels for all training images, which are combined with original labeled data to fine-tune the VLM. This iterative process of pseudo-label generation and optimization significantly enhances the VLM’s performance.
ZJYD CXY Vision proposes Instruction DINO (ISD), a DETR-based detector architecture and incorporates early fusion of image and text information, using a Swin-L visual backbone and EVA02-CLIP-L text encoder. Pre-training involves two stages using various datasets, transforming grounding data into single-object descriptions with QWen Max. For few-shot fine-tuning, the model adopts a flexible training format and uses VLMs like CLIP, TAP, and LLava for negative sample generation, finding that prompt tuning and text encoder fine-tuning generalize better than visual encoder fine-tuning. The final fine-tuning method combines prompt tuning and negative sampling, significantly improving mAP. To address sparse annotations, the visual encoder is initially fine-tuned to generate pseudo-label annotations, which are then used to complete training with prompt tuning.
Figure 5: Iteratively Prompting ChatGPT. Despite its large-scale pre-training, multi-modal models like ChatGPT-4o also suffers from concept alignment. Specifically, GPT-4o makes highly confident but incorrect predictions for debris. We propose an iterative prompting strategy to better align the model to a target concept. Given a few visual examples per-class from the training-set, we query ChatGPT to use its “web-scale knowledge” to generate text descriptions. We then augment the input to MQDet to incorporate this additional context for zero-shot evaluation.
Appendix C Iterative Prompting with Multi-Modal Chat Assistants
Typically, clients provide human annotators with a set of multi-modal instructions and a corpus of unlabeled data for annotation. Annotators start by first labeling a small subset of the data for review by the client, who acts as a domain expert and provides feedback on erroneous annotations (highlighting concept misalignment). Annotators use this feedback to annotate another subset of the data. This iterative process continues until the client is satisfied with the annotators’ ability to accurately label the entire dataset.
As shown in Figure 5, we explore the idea of iteratively prompting multi-modal chat assistants like ChatGPT to mimic the real-world workflow of human annotators. We start by asking GPT-4o to classify image crops of debris (derived from the few-shot training split). Notably, GPT-4o incorrectly classifies these training examples with high confidence. Therefore, we prompt GPT-4o to generate its own text descriptions of the few-shot examples according to its “web-scale knowledge”. Finally, we use the class names, generated text descriptions for debris, and few-shot visual examples to MQDet to predict instances of debris in the test-set.
We find that prompting MQDet with class names, ChatGPT generated text descriptions, and few-shot visual examples improves performance by 0.67% ($21.42$ mAP $\rightarrow$ $22.09$ mAP) over the baseline. Interestingly, although debris does not change when prompted with generated text descriptions, pushable pullable ($3.6$ AP $\rightarrow$ $15.29$ AP) and barrier ($11.6$ AP $\rightarrow$ $15.31$ AP) accuracy improve significantly. We posit that this improvement is due to the reduction in confusion (or the over-confident incorrect predictions) between debris and pushable-pullable (and barrier). Surprisingly, one of the top submissions to our CVPR challenge also use ChatGPT to generate meaningful text descriptions to improve detection concept alignment.
Appendix D Analysis of Federated Fine-Tuning
Prior works follow the $K$-shot dataset creation process established by [wang2020frustratingly]. Importantly, each image in the dataset is exhaustively annotated for a subset of all classes. Recall, a federated dataset is also comprised of images that are exhaustively annotated for a specific category. This suggests that we can leverage existing insights about federated datasets [gupta2019lvis, zhou2021probablistic] to train better few-shot object detectors.
Fine-Tuning with FedLoss. We fine-tune Detic with Federated Loss (FedLoss) [zhou2021probablistic] using a subset $S$ of classes $C$ for each training image. Specifically, we use a binary cross-entropy loss on all classes in $S$ and ignore classes outside of $S$ during training. $S$ is comprised of the ground-truth annotation class along with randomly sampled negative classes for each image. We sample these negative classes in proportion to their square-root frequency in the training set. We find that probablistically sampling negatives rather than labeling all unannotated classes as negatives improves fine-tuning results, reliably beating zero-shot performance. Importantly, although FedLoss has been explored in the context of long-tailed detection, applying it to FSOD provides considerable performance improvements, reaffirming that FSOD benchmarks are actually federated datasets.
Fine-Tuning with Pseudo-Negative Federated Loss (Ours). Despite the effectiveness of FedLoss, probablistically sampling negatives using dataset-wide statistics is sub-optimal because it does not consider the content of each image. We can improve the accuracy of sampled negatives with pseudo-labels to determine which classes are likely not in a particular image. If the maximal score for any class prediction is less than a threshold, we consider this class to be a negative. Using zero-shot model predictions to identify pseudo-negatives yields better results than simply using dataset-wide statistics. We find that this strategy works the best. We present pseudo-code in Alg. 1. All federated fine-tuning results in the main paper are trained with psuedo-negative federated loss.
@@ Appendix D Analysis of Federated Fine-Tuning
Approach 10 Shots: Average Precision (AP)
@@
Detic (Zero-Shot) [zhou2022detecting] 14.26 27.28 15.15 2.36
   + Standard Fine-Tuning 15.53 26.01 18.02 3.88
   w/ FedLoss 15.57 27.20 18.13 2.89
   w/ Pseudo-Negatives 16.67 29.15 18.71 3.90
   w/ True Negatives (Oracle) 16.99 29.60 18.94 4.21
   w/ Exhaustive Annotations (Oracle) 18.51 33.51 20.30 3.93
Oracle Performance Analysis. We empirically validate the effectiveness of our pseudo-negative federated loss by computing the upper bound performance when given access to ground-truth negatives and exhaustive annotations for the few-shot data split. Recall, nuImages is exhaustively annotated, but is repurposed for Foundational FSOD.
To compute the set of ground-truth negatives for each image, we use exhaustive ground-truth annotations to determine which categories are not present. Training with ground-truth negatives provides an upper bound on our pseudo-negatives experiment. Next, we train using exhaustive ground-truth annotations to provide an upper bound for the specific set of images used during training. In addition, this experiment highlights the performance gap between having exhaustive negatives and exhaustive annotations.
Despite strong performance on classes with many and medium, the upper bound for classes with few examples remains low (4.21 AP and 3.93 AP). Given the success of training with pseudo-negatives, a natural next-step is to train with pseudo-positives. Our preliminary results suggest that incorporating pseudo-positives does not provide significant improvement over simply training with pseudo-negatives. We posit that training with incorrect pseudo-positives may incur a higher penalty than training with incorrect pseudo-negatives. This is a promising direction for future work.
Appendix E Impact of Box-Level Supervision for Foundational FSOD
We evaluate the importance of using bounding-box supervised data in pre-training. Unlike Detic, which trains on box-supervised data from LVIS, COCO and image-text data from ImageNet21k, RegionCLIP[zhong2022regionclip] only pre-trains on image-text pairs from the Conceptual Captions (CC3M) dataset [cc4m].
@@ Appendix E Impact of Box-Level Supervision for Foundational FSOD
RegionCLIP (Zero-Shot) [zhong2022regionclip] 2.34 3.33 3.45 0.22
Detic (Zero-Shot) [zhou2022detecting] 14.26 27.28 15.15 2.36
RegionCLIP (Fine-Tuning, 5 shots) [zhong2022regionclip] 3.61 6.20 4.63 0.26
Detic (Fine-Tuning, 5 shots) [zhou2022detecting] 14.50 24.09 16.90 3.70
RegionCLIP (Fine-Tuning, 10 shots) [zhong2022regionclip] 3.58 6.10 4.65 0.24
Detic (Fine-Tuning, 10 shots) [zhou2022detecting] 15.28 26.93 18.00 3.27
RegionCLIP (Fine-Tuning, 30 shots) [zhong2022regionclip] 3.57 6.13 4.61 0.22
Detic (Fine-Tuning, 30 shots) [zhou2022detecting] 16.65 27.45 19.46 4.02
Table 9: Diagnosing RegionCLIP’s Poor Zero-Shot Performance. RegionCLIP’s zero-shot performance lags far behind Detic. Using RegionCLIP’s classifier on ground-truth region proposals yields high performance, suggesting that RegionCLIP struggles to accurately distinguish between foreground-vs-background.